# SPECIAL RELATIVITY

### v9

@@ 4.2 The ”Plane Wave” Emitter
We shall try to understand the effect in a very simple case. The explanation of the effect of aberration is actually based on the use of the model of single plane-wave emitter. As a simple model of a plane-wave emitter, we use a two-dimensional array of identical coherent elementary sources (dipoles), uniformly distributed on a given ($x-y$) plane $P$. We take the elementary sources to start radiating waves simultaneously with respect to the lab reference frame where the plane $P$ is at rest. Therefore we have a plane full of sources, oscillating together, with their motion in the plane and all having the same amplitude and phase. Let us suppose that the elementary sources are oscillating at frequency $\omega$. By letting the distance between each two adjacent elementary sources approach zero (i.e. much smaller with respect to the radiation wavelength $\lambda=2\pi c/\omega$), we may consider this two-dimensional arrangement as an ideal plane-wave emitter.
The concept of an (infinite) plane wave is widely used in physics. It is an analytically well-behaved solution of Maxwell’s equations. However, it is not a physically realizable solution, because the total energy content of such a wave is infinite. In any physically realizable situation, one will have to consider a finite source aperture. It is always hiddenly assumed that the detector for the direction of the radiation is an energy propagation detector and the size of the detector aperture is sufficiently large compared with the radiation beam size. Indeed, what is usually considered as an aberration is, in fact, an apparent deviation of the energy transport direction.
@@
There could be two approaches to the analysis of the aberration o light radiated by a single moving emitter. The first one is the covariant approach. The explanation of the effect of aberration of light in the case of a single moving emitter presented in the literature is actually based on the use of a Lorentz boost to describe how the direction of a beam of light depends on the velocity of the light source relative to the inertial frame of reference. Another non-covariant approach consists in using a ”single frame” description without reference to Lorentz transformations.
@@
In this section we demonstrate both approaches. Let us start with non-covariant approach. We must emphasize that there is no principle difficulty with the a non-covariant (3+1) approach in relativistic electrodynamics. It is perfectly satisfactory. The aberration of light problem can be treated within the same ”single inertial frame” description without reference to Lorentz transformations. Within a single inertial frame, the time is an absolute quantity in special relativity also. What does ”absolute” time mean? It means that simultaneity is absolute and there is no mixture of positions and time when sources change their velocities in the inertial frame. A distinctive trait of our non-covariant theory is the absence of relativistic kinematics in the description of aberration of light phenomena. When one has a transversely moving emitter there is the deviation of the energy transport for radiated light. According to the ”single frame” approach, this effect is a consequence of the fact that the Doppler effect is responsible for the angular frequency dispersion of the radiated light waves (Fig. 1 left).
Figure 1: The case when the $(x-y)$ plane emitter in an inertial lab frame is accelerated from rest up to velocity $v$ along the $x$-axis. It is assumed that there is no physical influence of the detector for the direction of the radiation on the measurement. The aberration increment is connected with the physical parameters by the relation: $\theta_{a}=v/c$. The explanation of the effect of aberration is based on the use of a Galilean (left) and Lorentz (right) boost to describe how the direction of a light beam depends on the velocity of light source relative to the lab frame. If make a Lorentz boost, we automatically introduce a time transformation $t\to t+xv/c^{2}$ and the effect of this transformation is just a rotation of the radiation phase front in the lab frame.
Let us consider the electrodynamics of the moving source. The explanation of the phenomenon of radiation in our case of interest consists in using a (active) Galileo boost to describe the uniform translational motion of the source in the inertial lab frame. Maxwell’s equations are not preserved in form by the Galilean transformation, i.e. Maxwell’s equations are not invariant under Galilean transformation. The new terms that have to be put into the field equations due to use of Galilean transformation lead to the prediction of the Doppler effect.
One should not to be surprised to find that electrodynamics problem of moving emitter has intrinsic anisotropy. If fact, anisontropy results directly from the time-dependence of the transverse position of the moving emitter with finite aperture. What must be recognized is that in the time-dependent emitter problem, the results will depend on the direction of the velocity vector.
The present approach to moving emitter problem uses the Fourier transform methods. When we are dealing with linear systems it is useful to decompose a complicated input into a number of more simple inputs, to calculate the response of the system to each of these elementary functions, and to superimpose the individual responses to find the total response. Fourier analysis provides the basic means for performing such a decomposition (1).
@@
The emitter with finite aperture is a kind of active medium which breaks up the radiated beam into a number of diffracted beams of plane waves. Each of these beams corresponds to one of the Fourier components into which an active medium can be resolved. Let us assume that the dipole density of the elementary source varies according to the law $\rho_{dip}=g(K_{\perp})\cos{({K}_{\perp}x)}$. We conclude, then, that the active medium of the elementary source is sinusoidally space-modulated.
@@
There is a different physical viewpoint on the Doppler effect of the light beam radiated from a moving emitter that is equivalent to the presented above. The phase of the wave at the world point $(\vec{r},t)$ is invariant at a change of the reference system. This invariance is independent of the coordinate transformation, by which we describe the change of the reference system. Therefore, the phase $\vec{k}\cdot\vec{r}-\omega t$ must be invariant of the Galilean transformation. Consequently, $\vec{k}^{\prime}\cdot\vec{r}^{\prime}-\omega^{\prime}t^{\prime}=\vec{k}\cdot\vec{r}-\omega t$. Substituting the Galilean transformation formulae $x^{\prime}=x-vt,t^{\prime}=t$ into the phase equality formula we obtain $\Delta\omega=K_{\perp}v$. This frequency change coincides with the result derived directly from the dispersion equation, as must be.
@@
Let us first remind the reader of the fact that the usual velocity of waves is defined as given the phase difference between the oscillations observed at two different points in a free plane wave. It is primary used for computing interference fringes that makes phase differences visible. In a plane wave we observe the phase velocity $\omega/k$. Another (group, or energy propagation) velocity can be defined, if we consider the propagation of a peculiarity, that is change in amplitude impressed on a train of waves. A simple combination of groups obtains when two waves $\omega_{1}=\omega+\Delta\omega,~{}k_{1}=k+\Delta k$ and $\omega_{2}=\omega-\Delta\omega,~{}k_{2}=k-\Delta k$ are superimposed. This represents a carrier with frequency $\omega$ and a modulation with frequency $\Delta\omega$. The wave may be described as a succession of moving beats (or groups). The carrier’s velocity is $\omega/k$, while the group velocity is given by $v_{g}=\Delta\omega/\Delta k\to d\omega/dk$.
@@
Many textbooks on electromagnetic theory discuss the aberration of light phenomena in the context of plane wave. However, in dealing with plane wave one will have an incorrect model of the aberration of light. When an infinite sinusoidal wave travels, there is a uniform average energy density throughout the space. Does this energy remain where it is, or does it propagate through the space? It is impossible to know this. All experimental methods for measuring the aberration of light operate with light signals, and hence do not measure the phase velocity but the signal velocity and this velocity coincides with group velocity.
@@
In our example, the plane waves with different wavenumber vectors propagate out from the moving emitter with the different frequencies. Then equation $\Delta\omega/\Delta k_{x}=v$ holds for each scattered waves independently on the sign and the magnitude of the radiated angle. In fact, the $\Delta\omega$ in our case of interest is the Doppler shift $\Delta\omega=\vec{K}_{\perp}\cdot\vec{v}$ and the $\Delta k_{x}$ is simply the transverse component of the radiated wavenumber vector $\Delta k_{x}=K_{\perp}$. The last equations state that radiated light beam with finite transverse size moves along the $x$ direction with group velocity $d\omega/dk_{x}=v$.
@@
It should be note, however, that there is another satisfactory way of explaining the effect of aberration of light from the moving source. The explanation consists in using a Lorentz boost to describe the uniform translation motion of the light source in the lab frame. On the one hand, the Maxwell’s equations remain invariant with respect to Lorentz transformations. On the other hand, if make a Lorentz boost, we automatically introduce a time transformation $t^{\prime}=t-xv/c^{2}$ and the effect of this transformation is just a rotation of the radiation phase front in the inertial frame. This is because the effect of this time transformation is just a dislocation in the timing of processes, which has the effect of rotating the plane of simultaneity on the angle $v/c$ in the first order approximation.
@@
According to Maxwell’s electrodynamics, coherent radiation is always emitted in the direction normal to the radiation phase front. This is because Maxwell’s equations have no intrinsic anisotropy. In other words, when a uniform translational motion of the source is treated according to Lorentz transformations, the aberration of light effect is described in the language of relativistic kinematics. According to the relativistic kinematics, the extra phase chirp $d\phi/dx=k_{x}=v\omega_{i}/c^{2}$ is introduced and the array of identical elementary sources of the moving emitter now have different phases. As a consequence of this, the plane wave wavefront rotates after the Lorentz transformation. Then, the radiated light beam is propagated at the angle $v/c$, yielding the phenomenon of the aberration of light (Fig. 1 right).
@@
We now ask about the group velocity of the radiated beam. With a plane wave $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ dispersion equation in the case of Maxwell’s electrodynamics is reduced to $k_{z}^{2}+k_{x}^{2}-\omega^{2}/c^{2}=0$. From the initial conditions and the Lorentz transformation we find that $\omega=(\omega_{i}+vK_{\perp})$, $k_{z}=\sqrt{\omega^{2}_{i}/c^{2}-K_{\perp}^{2}}$, $k_{x}=(v\omega_{i}/c^{2}+K_{\perp})$, where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated dipole density, $\omega_{i}$ is the frequency of the emitter radiation before the acceleration. Substituting these expressions in dispersion equation we find that the latter is satisfied, as must be.
@@
As one of the consequences of the Doppler effect in the Lorentz coordinatization, we find an angular frequency dispersion of the light waves radiated from the moving emitter with finite aperture. The Doppler shift, $\Delta\omega$, of radiated light wave (in the first order approximation) is given by $\Delta\omega=\vec{K}_{\perp}\cdot\vec{v}$, where $K_{\perp}$ is the transverse component of the radiated wavenumber vector. The last equation state that radiated light beam with finite transverse size moves along the $x$ direction with group velocity $d\omega/dk_{x}=v$.
@@
It is interesting to discuss what it means that there are two different (covariant and non-covariant) approaches that produce the same group velocity. The point is that both approaches describe correctly the same physical reality and since the group velocity has obviously an objective meaning (i.e. convention-invariant), both approaches yield the same physical results.
@@
The difference between the absolute time coordinatization and the Lorentz coordinatization is very interesting. In the Chapter 3 we already discussed how one can transform the absolute time coordinatization to Lorentz coordinatization. We can interpret manipulations with rule-clock structure in the lab frame simply as a change of the time variable according to the transformation $t\to t+xv/c^{2}$. The overall combination of Galileo transformation and variable changes actually yields the Lorentz transformation in the case of absolute time coordinatization in the inertial frame. This variable change has no intrinsic meaning. One can see the connection between the time shift and the issue of clock synchrony. The convention-independent results of calculations are precisely the same in the new variables. As a consequence, we should not care to transform the results of the electrodynamics problem solution into the original (3+1) variables.
@@
An idea of studying the relativistic electrodynamics using technique involving a change of variables is useful from a pedagogical point of view. It is worth remarking that the absent of a dynamical explanation for wavefront rotation in the Lorentz coordinatization has disturbed some physicists. It should be clear from the discussion in the Chapter 3 that a good way to think of the wavefront rotation is to regard it as a result of transformation to a new time variable in the framework of the Galilean (”single frame”) electrodynamics.
@@
4.5 Reflection from a Mirror Moving Transversely. Mistake in Existing Theory
@@
First, we examine the reasoning presented in textbooks (2). The reflection from the mirror is analyzed in two Lorentz reference frames. The fixed (lab) frame $S$ is at rest with respect to the plane-wave emitter. The moving frame $S^{\prime}$ has velocity $v$. In this frame, the mirror is at rest. In both frames, we use a Cartesian coordinate system in which $x-y$ plane is tangent to the reflection surface. The $x$ direction coincides with the direction of $v$. For simplicity, we consider the case in which light is incident from the $z$ direction in the lab frame. Incident light is described by its four-dimensional wave vector, whose time like component is the angular frequency $\omega$ and whose space like components define the direction of propagation. In the lab frame, $(t,x,y,z)$, this vector has components $k_{1}=(\omega,0,0,-\omega/c)$, where the negative sign indicates propagation towards the mirror (Fig. 3a). Our task is to determine the wave vector for the reflected beam.
@@
Figure 2: Transversely moving mirror with small aperture at normal incidence. According to textbooks, there is no deviation of the energy transport for the reflected light beam. A monochromatic plane wave of light is falling normally on the small moving aperture mirror, and generates a reflected oblique beam.
@@
Figure 3: The effect of aberration of light is described in the language of relativistic kinematics, in terms of the wavenumber four vector. Geometry of the reflection as seen from (a) lab frame, (b) inertial frame moving with the same velocity as the mirror. According to textbooks, there is no aberration for light reflected from a tangentially moving mirror.
@@
The argument that there is no aberration for light reflected from a transversely moving mirror runs something like this. It is easiest to consider the reflection in the moving frame (Fig. 3b). In this frame, the surface is at rest, so the usual laws of optical reflection apply. We will describe the effect of aberration of light by working only up to the first order $v/c$. An observer moving with the mirror surface sees the wave vector $k_{1}^{\prime}=(\omega,-v\omega/c^{2},0,-\omega/c)$. The effect of reflection is to reverse the sign of the $z^{\prime}$ component of the wave vector, $k^{\prime}_{2}=(\omega,-v\omega/c^{2},0,\omega/c)$. We now obtain the reflected wave vector in the lab frame by applying the inverse Lorentz transformation: $k_{2}=(\omega,0,0,\omega/c)$.This vector represents a light beam traveling away from the mirror, having the same frequency as the incoming beam. This shows that the beam is reflected according to the usual geometrical optics laws, and the beam suffers no aberration.
@@
There are two errors in solving problems concerning reflection from a mirror moving transversely. The first error in the last argument follows from the fact that the concept of a plane wave and an infinite plane mirror is used in the case of the tangential motion. The absurdity of this, i.e. the impossibility of such a motion ever been detected is put forward as an obvious argument against the reasoning presented in textbooks. Indeed, if the mirror is infinite the problem is not time dependent (we must conclude that when we are dealing with hidden assumption that problem is not time dependent we have no aberration and that is not too surprising). It is obvious that only the motion of a finite mirror is meaningful. In view of the electrodynamics, only a velocity of the finite mirror has a physical meaning. The reasoning in textbooks ignores completely the interaction of the light and the moving mirror edges.
We would also noticed that the concept of plane wave is used in textbooks. There is a common mistake made in the electrodynamics connected with the energy transport direction in the case of an (infinite) plane wave. When an infinite plane wave travels, there is a uniform average energy density throughout the space. It is impossible to know the energy transport direction when one has deal with a plane wave. All experimental methods for measuring the aberration increment operate with light signals, and hence do not measure the phase velocity (i.e. frequency and wavenumber vector) but the group velocity. It can be defined only if we consider the propagation of a peculiarity, that is change in amplitude impressed on a train of waves. What authors of textbooks generally overlooked is the fact that the energy transport problem is well-defined only if the source and mirror apertures have already been specified.
@@
The second error in the presented textbook argument is conceptual in nature. Suppose that an observer, which is at rest with respect to the inertial frame performs the direction of the energy transport measurement. The observer describes the light beam from the emitter at rest using diagonal (Minkowski) metric. According to textbooks, the metric of the moving mirror is the same (i.e. Minkowski metric) as metric of the emitter at rest. The Maxwell’s equations can be applied in the lab inertial frame only in the case when Lorentz coordinates are assigned. The authors of textbooks incorrectly belief that the common Lorentz time coordinate axis for emitter and mirror can be assigned. This is misconception. The question arises how to assign synchronization in the lab frame after the mirror acceleration.
@@
The peculiarity of this problem with the viewpoint of relativistic kinematics is that here the emitter (and observer with his measuring devices) is at rest in the lab inertial frame and the mirror is moving with the constant speed with respect to the lab frame and interacts with the radiated light beam. How can we solve a problem involving the emitter-mirror relative velocity? Suppose that we assign the Lorentz time coordinate for the description of the emitter radiation. But this will be the absolute time coordinatization for the boosted mirror and this boost will be described in such coordinatization by the Galilean transformation. Without changing synchronization in the lab frame we can prepare, for mirror and emitter, a common set of synchronized clocks only in the case of absolute time coordinatization i.e in the case when simultaneity is absolute.
@@
Suppose that we re-synchronize clocks in the lab frame in order to assign the Lorentz time coordinate for the boosted mirror. In this coordinatization, we describe the reflection of light using the usual Maxwell’s equations. On the other hand, the effect of this time transformation is just a rotation of the radiation phase front of the incoming plane wave on the angle $v/c$. As a result, this new time coordinate in the lab frame is interpreted by saying that Maxwell’s equations are not applicable to the emitter radiation description. It is not difficult to see that the peculiarity of this coordinatization is that here the energy transport velocity is not equal to the phase velocity for incoming light beam.

### v10

@@ 4.1 The ”Plane Wave” Emitter
We shall try to understand the effect in a very simple case. The explanation of the effect of aberration is based on the use of the model of a single plane-wave emitter. As a simple model of a plane-wave emitter, we use a two-dimensional array of identical coherent elementary sources (dipoles), uniformly distributed on a given ($x-y$) plane $P$. We take the elementary sources to start radiating waves simultaneously with respect to the lab reference frame where the plane $P$ is at rest. Therefore we have a plane full of sources, oscillating together, with their motion in the plane and all having the same amplitude and phase. Let us suppose that the elementary sources are oscillating at frequency $\omega$. By letting the distance between each two adjacent elementary sources approach zero (i.e. much smaller with respect to the radiation wavelength $\lambda=2\pi c/\omega$), we may consider this two-dimensional arrangement as an ideal plane-wave emitter.
The concept of an (infinite) plane wave is widely used in physics. It is an analytically well-behaved solution of Maxwell’s equations. However, it is not a physically realizable solution, because the total energy content of such a wave is infinite. In any physically realizable situation, one will have to consider a finite source aperture. It is always hiddenly assumed that the detector for the direction of the radiation is an energy propagation detector and the size of the detector aperture is sufficiently large compared with the radiation beam size. Indeed, what is usually considered an aberration is, in fact, a deviation of the energy transport direction.
@@
There could be two approaches to the analysis of the aberration of light radiated by a single moving emitter. The first one is the covariant approach. The explanation of the effect of aberration of light in the case of a single moving emitter presented in the literature is based on the use of a Lorentz boost to describe how the direction of a beam of light depends on the velocity of the light source relative to the inertial frame of reference. Another non-covariant approach consists of using a ”single frame” description without reference to Lorentz transformations.
@@
In this section, we demonstrate both approaches. Let us start with the non-covariant approach. We must emphasize that there is no principle difficulty with the non-covariant (3+1) approach in relativistic electrodynamics. It is perfectly satisfactory. The aberration of light problem can be treated within the same ”single inertial frame” description without reference to Lorentz transformations. Within a single inertial frame, the time is an absolute quantity in special relativity also. What does ”absolute” time mean? It means that simultaneity is absolute and there is no mixture of positions and time when sources change their velocities in the inertial frame. A distinctive trait of our non-covariant theory is the absence of relativistic kinematics in the description of the aberration of light phenomena. When one has a tangentially (to its surface) moving emitter there is the deviation of the energy transport for radiated light. According to the ”single frame” approach, this effect is a consequence of the fact that the Doppler effect is responsible for the angular frequency dispersion of the radiated light waves (Fig. 1 left).
Figure 1: The case when the plane wave emitter in an inertial lab frame is accelerated from rest up to velocity $v$ along the $x$-axis. The aberration increment is connected with the physical parameters by the relation: $\theta_{a}=v/c$. The explanation of the effect of aberration is based on the use of a Galilean (left) and Lorentz (right) boost.
Let us consider the electrodynamics of the moving source. The explanation of the phenomenon of radiation in our case of interest consists of using a (active) Galileo boost to describe the uniform translational motion of the source in the inertial lab frame. Maxwell’s equations are not preserved in form by the Galilean transformation, i.e. Maxwell’s equations are not invariant under Galilean transformation. The new terms that have to be put into the field equations due to the use of Galilean transformation lead to the prediction of the Doppler effect.
The present approach to the moving emitter problem uses the Fourier transform method. When we are dealing with linear systems it is useful to decompose a complicated input into a number of more simple inputs, to calculate the response of the system to each of these elementary functions, and to superimpose the individual responses to find the total response. Fourier analysis provides the basic means for performing such a decomposition  [^16] For a general discussion of the Fourier transform methods of spatial filtering theory or Abbe diffraction theory we suggest reading the book [22]..
@@
The emitter with a finite aperture is a kind of active medium that breaks up the radiated beam into a number of diffracted beams of plane waves. Each of these beams corresponds to one of the Fourier components into which an active medium can be resolved. Let us assume that the dipole density of the elementary source varies according to the law $\rho_{dip}=g(K_{\perp})\cos{({K}_{\perp}x)}$. We conclude, then, that the active medium of the elementary source is sinusoidally space-modulated.
@@
There is a different physical viewpoint on the Doppler effect of the light beam radiated from a moving emitter that is equivalent to the one presented above. The phase of the wave at the world point $(\vec{r},t)$ is invariant at a change of the reference system. This invariance is independent of the coordinate transformation, by which we describe the change of the reference system. Therefore, the phase $\vec{k}\cdot\vec{r}-\omega t$ must be invariant of the Galilean transformation. Consequently, $\vec{k}^{\prime}\cdot\vec{r}^{\prime}-\omega^{\prime}t^{\prime}=\vec{k}\cdot\vec{r}-\omega t$. Substituting the Galilean transformation formulae $x^{\prime}=x-vt,t^{\prime}=t$ into the phase equality formula we obtain $\Delta\omega=K_{\perp}v$. This frequency change coincides with the result derived directly from the dispersion equation, as must be.
@@
Let us first remind the reader of the fact that the usual velocity of waves is defined as the phase difference between the oscillations observed at two different points in a free plane wave. It is primarily used for computing interference fringes that make phase differences visible. In a plane wave, we observe the phase velocity $\omega/k$. Another (group, or energy propagation) velocity can be defined, if we consider the propagation of a peculiarity, that is change in amplitude impressed on a train of waves. A simple combination of groups is obtained when two waves $\omega_{1}=\omega+\Delta\omega,~{}k_{1}=k+\Delta k$ and $\omega_{2}=\omega-\Delta\omega,~{}k_{2}=k-\Delta k$ are superimposed. This represents a carrier with frequency $\omega$ and a modulation with frequency $\Delta\omega$. The wave may be described as a succession of moving beats (or groups). The carrier’s velocity is $\omega/k$, while the group velocity is given by $v_{g}=\Delta\omega/\Delta k\to d\omega/dk$.
@@
Many textbooks on electromagnetic theory discuss the aberration of light phenomena in the context of plane wave. However, in dealing with plane waves one will have an incorrect model of the aberration of light. When an infinite sinusoidal wave travels, there is a uniform average energy density throughout the space. Does this energy remain where it is, or does it propagate through the space? It is impossible to know this. All experimental methods for measuring the aberration of light operate with light signals, and hence do not measure the phase velocity but the signal velocity and this velocity (in our case of interest) coincides with the group velocity.
@@
In our example, the plane waves with different wavenumber vectors propagate out from the moving emitter with different frequencies. Then equation $\Delta\omega/\Delta k_{x}=v$ holds for each scattered wave independently on the sign and the magnitude of the radiated angle. In fact, the $\Delta\omega$ in our case of interest is the Doppler shift $\Delta\omega=\vec{K}_{\perp}\cdot\vec{v}$ and the $\Delta k_{x}$ is simply the transverse component of the radiated wavenumber vector $\Delta k_{x}=K_{\perp}$. The last equations state that a radiated light beam with finite transverse size moves along the $x$ direction with group velocity $d\omega/dk_{x}=v$.
@@
It should be noted, however, that there is another satisfactory way of explaining the effect of aberration of light from the moving source. The explanation consists of using a Lorentz boost to describe the uniform translation motion of the light source in the lab frame. On the one hand, Maxwell’s equations remain invariant with respect to Lorentz transformations. On the other hand, if make a Lorentz boost, we automatically introduce a time transformation $t^{\prime}=t-xv/c^{2}$, and the effect of this transformation is just a rotation of the radiation phase front in the inertial frame. This is because the effect of this time transformation is just a dislocation in the timing of processes, which has the effect of rotating the plane of simultaneity on the angle $v/c$ in the first-order approximation.
@@
According to Maxwell’s electrodynamics, coherent radiation is always emitted in the direction normal to the radiation phase front. This is because Maxwell’s equations have no intrinsic anisotropy. In other words, when a uniform translational motion of the source is treated according to Lorentz transformations, the aberration of light effect is described in the language of relativistic kinematics. According to the relativistic kinematics, the extra phase chirp $d\phi/dx=k_{x}=v\omega_{i}/c^{2}$ is introduced and the array of identical elementary sources of the moving emitter now has different phases. As a consequence of this, the plane wave wavefront rotates after the Lorentz transformation. Then, the radiated light beam is propagated at the angle $v/c$, yielding the phenomenon of the aberration of light (Fig. 1 right).
@@
We now ask about the group velocity of the radiated beam. With a plane wave $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ dispersion equation in the case of Maxwell’s electrodynamics is reduced to $k_{z}^{2}+k_{x}^{2}-\omega^{2}/c^{2}=0$. From the initial conditions and the Lorentz transformation we find that $\omega=(\omega_{i}+vK_{\perp})$, $k_{z}=\sqrt{\omega^{2}_{i}/c^{2}-K_{\perp}^{2}}$, $k_{x}=(v\omega_{i}/c^{2}+K_{\perp})$, where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated dipole density, $\omega_{i}$ is the frequency of the emitter radiation before the acceleration. Substituting these expressions in the dispersion equation we find that the latter is satisfied, as must be.
@@
As one of the consequences of the Doppler effect in the Lorentz coordinatization, we find an angular frequency dispersion of the light waves radiated from the moving emitter with a finite aperture. The Doppler shift, $\Delta\omega$, of a radiated light wave (in the first order approximation) is given by $\Delta\omega=\vec{K}_{\perp}\cdot\vec{v}$, where $K_{\perp}$ is the transverse component of the radiated wavenumber vector. The last equation states that a radiated light beam with finite transverse size moves along the $x$ direction with group velocity $d\omega/dk_{x}=v$.
@@
It is interesting to discuss what it means that there are two different (covariant and non-covariant) approaches that produce the same group velocity. The point is that both approaches describe correctly the same physical reality and since the group velocity has an objective meaning (i.e. convention-invariant), both approaches yield the same physical results.
@@
The difference between the absolute time coordinatization and the Lorentz coordinatization is very interesting. In Chapter 3 we already discussed how one can transform the absolute time coordinatization to Lorentz coordinatization. We can interpret manipulations with a rule-clock structure in the lab frame simply as a change of the time variable according to the transformation $t\to t+xv/c^{2}$. The overall combination of Galileo transformation and variable changes actually yields the Lorentz transformation in the case of absolute time coordinatization in the inertial frame. This variable change has no intrinsic meaning. One can see the connection between the time shift and the issue of clock synchrony. The convention-independent results of calculations are precisely the same in the new variables. As a consequence, we should not care to transform the results of the electrodynamics problem solution into the original (3+1) variables.
@@
The idea of studying relativistic electrodynamics using a technique involving a change of variables is useful from a pedagogical point of view. It is worth remarking that the absence of a dynamical explanation for wavefront rotation in the Lorentz coordinatization has disturbed some physicists. It should be clear from the discussion in Chapter 3 that a good way to think of the wavefront rotation is to regard it as a result of the transformation to a new time variable in the framework of the Galilean (”single frame”) electrodynamics.
@@
4.4 Reflection from a Mirror Moving Parallel to Itself. Mistake in Existing Theory
@@
First, we examine the reasoning presented in textbooks  [^17] It is generally believed that there is no aberration for light reflected from mirrors moving parallel to itself. To quote e.g. Sommerfeld [23]: ”Thus, for a mirror moving tangentially to its surface the law of reflection which holds for the stationary mirror is preserved.” Similar statements can also be found in other textbooks. To quote e.g. Ugarov [24]: ” Hence, when the mirror moves parallel to itself the frequency of incident light is equal to that of reflected light, and the angle of incidence is equal to the angle of reflection.”. The reflection from the mirror is analyzed in two Lorentz reference frames. The fixed (lab) frame $S$ is at rest with respect to the plane-wave emitter. The moving frame $S^{\prime}$ has velocity $v$. In this frame, the mirror is at rest. In both frames, we use a Cartesian coordinate system in which the $x-y$ plane is tangent to the reflection surface. The $x$ direction coincides with the direction of $v$. For simplicity, we consider the case in which light is incident from the $z$ direction in the lab frame. Incident light is described by its four-dimensional wave vector, whose time-like component is the angular frequency $\omega$ and whose space-like components define the direction of propagation. In the lab frame, $(t,x,y,z)$, this vector has components $k_{1}=(\omega,0,0,-\omega/c)$, where the negative sign indicates propagation towards the mirror (Fig. 3a). Our task is to determine the wave vector for the reflected beam.
@@
Figure 2: A mirror moves parallel to itself. According to textbooks, there is no deviation of the energy transport for the reflected light beam. A monochromatic plane wave of light falls normally on the small moving aperture mirror and generates a reflected oblique beam.
@@
Figure 3: The effect of aberration of light is described in the language of relativistic kinematics, in terms of the wavenumber four-vector. Geometry of the reflection as seen from (a) lab frame, (b) inertial frame moving with the same velocity as the mirror. According to textbooks, there is no aberration for light reflected from a tangentially moving mirror.
@@
The argument that there is no aberration for light reflected from a tangentially moving mirror runs something like this. It is easiest to consider the reflection in the moving frame (Fig. 3b). In this frame, the surface is at rest, so the usual laws of optical reflection apply. We will describe the effect of aberration of light by working only up to the first order $v/c$. An observer moving with the mirror surface sees the wave vector $k_{1}^{\prime}=(\omega,-v\omega/c^{2},0,-\omega/c)$. The effect of reflection is to reverse the sign of the $z^{\prime}$ component of the wave vector, $k^{\prime}_{2}=(\omega,-v\omega/c^{2},0,\omega/c)$. We now obtain the reflected wave vector in the lab frame by applying the inverse Lorentz transformation: $k_{2}=(\omega,0,0,\omega/c)$. This vector represents a light beam traveling away from the mirror, having the same frequency as the incoming beam. This shows that the beam is reflected according to the usual geometrical optics laws, and the beam suffers no aberration.
@@
There are two errors in solving problems concerning reflection from a mirror moving tangentially to its surface. The first error in the last argument follows from the fact that the concept of a plane wave and an infinite plane mirror is used in the case of tangential motion. The absurdity of this, i.e. the impossibility of such a motion ever being detected is put forward as an obvious argument against the reasoning presented in textbooks. Indeed, if the mirror is infinite the problem is not time-dependent (we must conclude that when we are dealing with the hidden assumption that the problem is not time-dependent we have no aberration and that is not too surprising). Only the motion of a finite mirror is meaningful. In view of the electrodynamics, only the velocity of the finite mirror has a physical meaning. The reasoning in textbooks ignores completely the interaction of the light and the moving mirror edges. What authors of textbooks generally overlooked is the fact that the energy transport problem is well-defined only if the source and mirror apertures have already been specified.
@@
The second error in the presented textbook argument is conceptual in nature. Suppose that an observer, which is at rest with respect to the inertial frame performs the direction of the energy transport measurement. The observer describes the light beam from the emitter at rest using the diagonal (Minkowski) metric. According to textbooks, the metric of the moving mirror is the same (i.e. Minkowski metric) as the metric of the emitter at rest. Maxwell’s equations can be applied in the lab inertial frame only in the case when Lorentz coordinates are assigned. The authors of textbooks incorrectly believe that the common Lorentz time coordinate axis for the emitter and mirror can be assigned. This is a misconception. The question arises of how to assign synchronization in the lab frame after the mirror acceleration.
@@
The peculiarity of this problem with the viewpoint of relativistic kinematics is that here the emitter (and observer with his measuring devices) is at rest in the lab inertial frame and the mirror is moving with the constant speed with respect to the lab frame and interacts with the radiated light beam. How can we solve a problem involving the emitter-mirror relative velocity? Suppose that we assign the Lorentz time coordinate for the description of the emitter radiation. But this will be the absolute time coordinatization for the boosted mirror and this boost will be described in such coordinatization by the Galilean transformation. Without changing synchronization in the lab frame we can prepare, for mirror and emitter, a common set of synchronized clocks only in the case of absolute time coordinatization i.e. in the case when simultaneity is absolute.
@@
Suppose we re-synchronize clocks in the lab frame to assign the Lorentz time coordinate for the boosted mirror. In this coordinatization, we describe the reflection of light using Maxwell’s equations. On the other hand, the effect of this time transformation is just a rotation of the radiation phase front of the incoming plane wave on the angle $v/c$. As a result, this new time coordinate in the lab frame is interpreted by saying that Maxwell’s equations are not applicable to the emitter radiation description. It is not difficult to see that the peculiarity of this coordinatization is that here the energy transport velocity is not equal to the phase velocity for the incoming light beam (see footnote in Section 3.6).